A student fires a cannonball vertically upwards. The cannonball returns to the ground after a 6.00s flight. Determine all unknowns and answer the following questions. Neglect drag and the initial height and horizontal motion of the cannonball. Use regular metric units (ie. meters).


1
Expert's answer
2021-10-04T10:06:00-0400

Given:

"t=6\:\rm s"

"g=9.8\:\rm m/s^2"


Unknown:

"v_0,\; h_{\max}."


The time of motion to the up

"t^*=t/2=3\:\rm s"

The initial velocity

"v_0=gt^*=9.8*3=29.4\:\rm m/s"

The maximum height

"h_{\max}=\frac{v_0^2}{2g}=\frac{29.4^2}{2*9.8}=44.1\:\rm m"